OpenStudy (anonymous):

Enter the quadratic equatuion in standard form. 2x2-3= -4x-1

OpenStudy (saifoo.khan):

2x^2 + 4x -2 = 0

OpenStudy (anonymous):

2x^2 + 4x -2 = 0

OpenStudy (anonymous):

2x^2+4x-2=0

OpenStudy (saifoo.khan):

Copy Cats! :O